1.

A company that makes shopping carts for supermarkets and other stores recently purchased some new equipment that reduces the labor content of the jobs needed to produce the shopping carts. Prior to buying the new equipment, the company used 6 workers, who together produced an average of 90 carts per hour. Workers receive $16 per hour, and machine cost was $30 per hour. With the new equipment, it was possible to transfer one of the workers to another department, and equipment cost increased by $10 per hour while output increased by 5 carts per hour. a. Compute labor productivity under each system. Use carts per worker per hour as the measure of labor productivity. (Round your answers to 3 decimal places.) Before 15 carts per worker per hour 19 carts per worker per hour After b. Compute the multifactor productivity under each system. Use carts per dollar cost (labor plus equipment) as the measure. (Round your answers to 3 decimal places.) Before After carts/dollar cost carts/dollar cost C. Comment on the changes in productivity according to the two measures. Round your intermediate calculations to 3 decimal places and final answers to 2 decimal places. Labor productivity increased *b Multifactor productivity (click to select) by

2.

A manager checked production records and found that a worker produced 190 units while working 50 hours. In the previous week, the same worker produced 128 units while working 40 hours. a. Compute Current period productivity and Previous period productivity. Round your answers to 2 decimal places.) Current period productivity Previous period productivity Units / hr Units/ hr b. Did the workers productivity increase, decrease, or remain the same? (Round your intermediate calculations and final answer to 2 decimal places. Omit the % sign in your response.) Workers productivity by 7% (Click to select) decreased remain the same increased
Answer #1

1 Before new equipment

No. of workers=6

Average output=90 carts per hour

Labor productivity L1=Average output/Number of workers=90/6=15 carts per hour per worker.

After new equipment

No. of workers=5 (one transferred )

Average output=95 carts per hour

Labor productivity L2=Average output/Number of workers=95/5=19 carts per hour per worker.

b-

Before new equipment

Workers rate = $16 per worker per hour

Machine cost = $30 per hour

Average output = 90 carts per hour

Multifactor productivity=Average output/{(No of workers*Workers rate)+machine cost}

M1=90/{(6*16)+30}=0.71 carts per dollar cost

After new equipment

Workers rate = $16 per worker per hour

Machine cost = $40 per hour

Average output = 95 carts per hour

Multifactor productivity=Average output/{(No of workers*Workers rate)+machine cost}

M2=95/{(5*16)+40}=0.79 carts per dollar cost

c. Increase in labor productivity=(L2-L1)/L1=(19-15)/15=0.26 or 26 %

Increase in multifactor productivity=(M2-M1)/M1=(0.79-0.71)/0.71=0.11 or 11%

2. Previous productivity =Produced unit/Hours=190/50=3.8 units per hour

Current productivity= Produced unit/Hours=128/40=3.2 units per hour

clearly productivity is decrease

percentage decreased=(3.2-3.8)/3.8= -0.15 or decreased by 15%.